### v3

@@ 1 Introduction
Instruction-pretrained models (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022) such as large language models (LLMs) are able to interpret input strings (prompts) as instructions to carry out a task. Given the lack of a feedback cycle (e.g., back-propagation), this is known as in-context learning (ICL; Wei et al. 2022b; Okawa et al. 2023; Ganguli et al. 2022; Brown et al. 2020; to name a few). ICL has led their deployment in multiple product areas, such as writing assistance, in spite of their sensitivity to the prompt’s phrasing (Khatun and Brown, 2023; Xiong et al., 2024; Lu et al., 2022). In turn, this spawned a considerable body of work to determine optimal prompting techniques and wordings, which itself has been shown to be automatable (Zhou et al., 2023; Shin et al., 2020), and effective when the model has some sort of guidance (Wei et al., 2022c; Zhang et al., 2023; Richards, 2023).
@@ 2.1 Why Category Theory?
Many aspects of ML, and LLMs in particular, rely on solid, well-understood foundations (for example, gradient descent), but it is not entirely clear how these foundations come together (Shiebler et al., 2021). For LLMs in particular, their size, sensitivity to the input, sometimes opaque design, and inconsistent (random) outputs [^1] Randomness in the output may be desirable for more diverse content, and hence better user experience. We do not consider it a flaw. means that mathematical modelling of these models is difficult. Additionally, a single LLM may have multiple distinct downstream applications with carefully-tuned task-specific prompts. Hence, developing the language required to model LLM behavior must be general enough to factor in their complexity, but also sufficiently expressive to account for their adaptability. Category theory (Eilenberg and MacLane, 1945) is a branch of mathematics that aims to study analogies between different concepts in mathematics by abstracting out certain details, but retaining rigor. [^2] An apt analogy is that of Spivak (2014) (p. 376), who compares it to studying the game of tennis without studying the physics of every particle in the system. This makes it suitable as a language to describe LLMs.
@@ 2.2 Category Theory
Given that category theory utilizes many specific concepts and variable terminology and notation, we begin by reviewing them. For a detailed exposition of these concepts, see Mac Lane (1998) and Riehl (2016). We include examples wherever suitable to facilitate an introduction to this subject.
@@ 2.2.2 Diagrams, Functors, and Natural Transformations
Functors can be thought of as analogies between categories (Baez, 2006; Riehl, 2016). More importantly, they can also be transformed between one another through a natural transformation $\alpha\colon F\implies G$. It is normally said that natural transformations are where the real power of category theory lies (Mac Lane, 1998).
@@ 2.2.3 Structure and Classification of Categories
[…]
@@
If for all $X\in\mathbf{C}$ there is an arrow, unique up to isomorphism, that maps to some object $1$, $!\colon X\rightarrow 1$, this object is said to be terminal.
@@ 2.3 Further Assumptions
[…]
@@ 2.4.2 Assumptions on Prompt Execution
We assume the LLM to be deterministic, that it has followed the instructions correctly, and that its output is grammatical. When the user input is not conducive to do so (e.g., “multiply {X} by $3$”, where X is an extract of Finnegans Wake), we assume the LLM responds with a boilerplate string (e.g., “As a LLM I cannot perform this task”). All of these assumptions are only for simplicity: we can always consider the codomain the entire $\Sigma^{k}$, or map any failed outputs to a special token.
@@ 2.4.3 A Mathematical Model of Prompting and Prompt-Sensitivity
Two prompts $p,q$ describing the same task (e.g., $p:=\;$“multiply {X} by 3” and $q:=\;$“add {X} 3 times”) do not have guaranteed equal behavior, as LLMs are prompt-sensitive. We consider isomorphism to be over meaning (i.e., two paraphrases without loss of meaning are isomorphic). Hence $p$ and $q$ have distinct codomains and so $\tilde{p}\neq\tilde{q}$. Prompt application is associative up to isomorphism: suppose $p:=``\text{Expand }\{X\}$”, $q:=``\text{Get keypoints }\{Y\}$”, and $r:=$“Translate to West Frisian $\{Z\}$”. Then $(pq)r\cong p(qr)\cong pqr$.
@@ 3.3 An Example
It may be helpful to illustrate this with an example. Let Summ be a task-category whose objects are passages; and the arrows are prompts that take passages to their summaries, such as $f(X):=\;$“Summarize {X}”. Since it is possible to obtain a summary of a summary, the arrows are composable and associative, and hence Summ is a category. We assign the internal hom to be the string descriptions of the prompts.
Let Expand be the task-category with key points for writing passages, and passages, as objects. Its arrows are prompts that take key points to passages, such as $f(X):=\;$“Write a passage using the following: $\{X\}$”. Composition and associativity are symmetric to Summ: all expanded key points are new key points.
[…]
@@ 6.1 Prompting to Prompt
Much existing work in prompting can be construed as meta-prompting approaches, although called by different names. Below we limit ourselves to discussing strategies that fit this criterion. For an in-depth survey of prompting methods, see Liu et al. (2023) and Qiao et al. (2023).
@@
For pretraining perhaps the clearest example is instruction pretraining (Wei et al., 2022a; Sanh et al., 2022; Ouyang et al., 2022), where a generator provides synthetic tasks (prompts and context) to the model being trained. Self-instruct (Wang et al., 2023) is an efficient-yet-noisy algorithm designed for instruction pre-training that takes in a small set of prompts and generates new prompts. LM-BFF (Gao et al., 2021) is an approach for pre-training and fine-tuning language models via automated prompt generation. The authors show it is more effective than basic fine-tuning in various tasks.
For performance improvement, Shin et al. (2020) showed with AutoPrompt that generating prompts is more effective than fine-tuning and manual prompts in low-resource scenarios. They also showed this method lowers hallucination rates in models. More recently, Zhou et al. (2023) showed that it is possible to automate some of the prompt engineering work.
Meta-prompting approaches have also found success at lowering biases (Guo et al., 2022), event argument extraction (Dong et al., 2022), and jailbreak detection (Pryzant et al., 2023). Liu et al. (2022) showed that generating knowledge automatically for question answering (QA) improves downstream performance. For reasoning, Auto-CoT (Zhang et al., 2023) propose building the pathways for reasoning in chain-of-thought prompting (Wei et al., 2022c; Kojima et al., 2022) automatically. Closer to pure meta-prompting automation, Yao et al. (2023) introduced a prompting paradigm designed for reasoning, which uses a feedback loop to generate reasoning steps. This is effective when integrating external-world grounding. LAMBADA (Kazemi et al., 2023) is a backward-chaining algorithm that focuses on reasoning by extracting propositions and performing proving via chaining. Auto-GPT (Richards, 2023) is a tool designed to iteratively solve problems by generating its own prompts and including a database for memory.
@@ 6.2 Category Theory
Our work is the first, to our knowledge, to formally describe applications of LLMs with category theory. However, there is a broad push from the community to apply this field to machine learning (Shiebler et al., 2021), such as modelling gradient-based learning (Cruttwell et al., 2022; Elliott, 2018; Fong et al., 2019), aspects of natural language processing like semantics (Kartsaklis et al., 2013; Piedeleu et al., 2015), QA (de Felice et al., 2019), discourse analysis (Coecke et al., 2018), and negation (Shaikh et al., 2021). These usually rely on a finer-grained approach to model language involving a functor between grammars and semantics (Coecke et al., 2010). The application of category theory to computation is well-known: for example, via Turing categories (Longo and Moggi, 1990; Cockett and Hofstra, 2008). Turing categories are considerably more general than task-categories, and are more suitable to study computation as opposed to LLM behavior on specific tasks.
For the reader interested in learning more about category theory, we encourage them to read the works by Fong and Spivak (2019) and Spivak (2014) in the context of applied category theory; and Mac Lane (1998), Riehl (2016) and Awodey (2010) for more in-depth content.
@@ 7 Limitations
Strictly speaking, our framework is not fully realistic. First, computable functions are partial functions, and hence so are prompts. We do not account or evaluate this. We have worked with tokens as opposed to integers. This is not a problem since for every tokenized string there exists an equivalent vector representation, and hence a correspondence with FinVect, the category of vector spaces over a finite field. Our model of isomorphism is paraphrasing: stricter definitions exist, but not without their drawbacks (Rainer et al., 2012).
@@
Finally, we have not modeled stochasticity, and assumed that the LLM follows the instructions. This is not necessarily the case. It is possible to model probability with a Markov category (Fritz et al., 2023) such as Stoch. Markov categories are symmetric; but Prompt is not required to be.
@@ 8 Conclusion
Our work with Prompt showed that it is possible to abstract out common LLM ICL issues such as prompt sensitivity, generalizability, and user interaction; and still be able to model LLM interaction. We plan to extend our framework into a more detailed treatment of Prompt to account for stochasticity, enrichment, and the FinVect category, and link that to user preferences. Some prompting techniques, such as chain-of-thought, and its hybridization with Markov chains, such as Deepseek R1’s DeepSeek-AI (2025) can also be modeled with our framework, although we leave it for future work.
@@ Appendix A Proof of Lemma 1
It then follows that any morphism in $\text{Hom}_{\text{{Task}}}(-,-)$ [^6] The notation $f(-)$ is standard in category theory. It means “insert your object here.” will have (by construction) a corresponding exponential object, and hence that any natural-language description will be an element of Task.
@@ Appendix C Proof of
Although in category theory it is often sufficient to prove statements with diagrams and arguing that they commute (“proof by diagram”), in this section we provide a clearer proof of . For convenience we restate it below, and remind the reader of the properties of the hom functor. Since Prompt is right-closed, the internal hom $\hom(-,-)$, a bifunctor, mapping objects and arrows from ${}^{\text{op}}\times\text{{Prompt}}{}$ Prompt to Prompt. In other words, for objects $(Y,A)\in\text{{Prompt}}{}^{\text{op}}\times\text{{Prompt}}{}$ and arrows $f\colon X\rightarrow Y$, $g\colon B\rightarrow A$, for the internal hom $\text{hom}(f,g)$ the diagram below commutes
@@ D.2 Ideation
Ideation takes in a text and suggest ideas on how to improve it. The underlying task-category Idea has as objects passages, and as arrows ways to improve the (same) passage. Idea is a restricted version of Rewrite (only creative rewrites are allowed). Each object $X\in\text{{Idea}}$ has general-purpose arrows such as “Make it more concise”, “Make it longer”, and “Explain this to a 5 year old”. These three are our baseline prompts for this task. We assume that our baseline arrows belong to every exponential object $Z^{X}\in\text{{Idea}}$, but that they also have some disjoint, context-specific arrows: for example, if the context is “doing something for a long time can be boring, having hobbies and doing sports can refresh our mind”, an appropriate morphism would be “Add some details to support your claim that hobbies and sports can refresh our mind.” For our dataset we sampled 300 points from WikiText-103 (Merity et al., 2017) and the TOEFL11 corpus (Blanchard et al., 2013). The full prompt is in Prompt 2.
@@ D.3 Creativity
The underlying task-category Creat has as objects passages, and as arrows prompts that generate said linguistic-and-semantically natural transitions between them. Same as before, we assume that the exponential objects of Creat have disjoint elements. Our baseline prompts are “Write a paragraph to connect the left text and right texts,” “Insert a passage connecting the two passages”, and the empty string (implicitly passage completion). For this task we sampled 300 points of TOEFL11, and version 3.0.0 of the DailyMail/CNN corpus (Hermann et al., 2015). The full prompt is in Prompt 3.

### v4

@@ 1 Introduction
Instruction-pretrained models (wei2022finetuned; sanh2022multitask; NEURIPS2022_b1efde53) such as large language models (LLMs) are able to interpret input strings (prompts) as instructions to carry out a task. Given the lack of a feedback cycle (e.g., back-propagation), this is known as in-context learning (ICL; wei2022emergent; NEURIPS2023_9d0f188c; 10.1145/3531146.3533229; GPT3; to name a few). ICL has led their deployment in multiple product areas, such as writing assistance, in spite of their sensitivity to the prompt’s phrasing (Reliability; xiong2023llms; lu-etal-2022-fantastically). In turn, this spawned a considerable body of work to determine optimal prompting techniques and wordings, which itself has been shown to be automatable (zhou2023large; shin-etal-2020-autoprompt), and effective when the model has some sort of guidance (ChainOfThought; zhang2023automatic; AutoGPT).
@@ 2.1 Why Category Theory?
Many aspects of ML, and LLMs in particular, rely on solid, well-understood foundations (for example, gradient descent), but it is not entirely clear how these foundations come together (CatMLSurvey). For LLMs in particular, their size, sensitivity to the input, sometimes opaque design, and inconsistent (random) outputs [^1] Randomness in the output may be desirable for more diverse content, and hence better user experience. We do not consider it a flaw. means that mathematical modelling of these models is difficult. Additionally, a single LLM may have multiple distinct downstream applications with carefully-tuned task-specific prompts. Hence, developing the language required to model LLM behavior must be general enough to factor in their complexity, but also sufficiently expressive to account for their adaptability. Category theory (CategoryTheory) is a branch of mathematics that aims to study analogies between different concepts in mathematics by abstracting out certain details, but retaining rigor. [^2] An apt analogy is that of Spivak2 (p. 376), who compares it to studying the game of tennis without studying the physics of every particle in the system. This makes it suitable as a language to describe LLMs.
@@ 2.2 Category Theory
Given that category theory utilizes many specific concepts and variable terminology and notation, we begin by reviewing them. For a detailed exposition of these concepts, see MacLane and Riehl. We include examples wherever suitable to facilitate an introduction to this subject.
@@ 2.2.2 Diagrams, Functors, and Natural Transformations
Functors can be thought of as analogies between categories (Baez2; Riehl). More importantly, they can also be transformed between one another through a natural transformation $\alpha\colon F\implies G$. It is normally said that natural transformations are where the real power of category theory lies (MacLane).
@@ 2.2.3 Structure and Classification of Categories
  • An additional (bi)functor, known as the tensor product, $\otimes\colon\text{{C}}\times\text{{C}}\rightarrow\text{{C}}$.
@@
If for all $X\in\mathbf{C}$ there is an arrow, unique up to isomorphism, that maps to some object $1$, $\colon X\rightarrow 1$, this object is said to be terminal.
@@ 2.3 Further Assumptions
Throughout this paper we will describe morphisms based on string descriptions as (e.g.) $f(x)=x^{2}:=\text{``Take the square of \{x\}''}$. For simplicity, we refer to $Z^{X}$ as an exponential object, even when $\otimes$ is not the Cartesian product. All our categories are locally small.
@@ 2.4.2 Assumptions on Prompt Execution
We assume the LLM to be deterministic, that it has followed the instructions correctly, and that its output is grammatical. When the user input is not conducive to do so (e.g., “multiply {X} by 33”, where X is an extract of Finnegans Wake), we assume the LLM responds with a boilerplate string (e.g., “As a LLM I cannot perform this task”). All of these assumptions are only for simplicity: we can always consider the codomain the entire $\Sigma^{k}$, or map any failed outputs to a special token.
@@ 2.4.3 A Mathematical Model of Prompting and Prompt-Sensitivity
Two prompts $p,q$ describing the same task (e.g., p:=p:=\;“multiply {X} by 3” and q:=q:=\;“add {X} 3 times”) do not have guaranteed equal behavior, as LLMs are prompt-sensitive. We consider isomorphism to be over meaning (i.e., two paraphrases without loss of meaning are isomorphic). Hence $p$ and $q$ have distinct codomains and so $\tilde{p}\neq\tilde{q}$. Prompt application is associative up to isomorphism: suppose $p:=``\text{Expand }\{X\}$”, $q:=``\text{Get keypoints }\{Y\}$”, and r:=r:=“Translate to West Frisian $\{Z\}$”. Then $(pq)r\cong p(qr)\cong pqr$.
@@ 3.3 An Example
It may be helpful to illustrate this with an example. Let Summ be a task-category whose objects are passages; and the arrows are prompts that take passages to their summaries, such as f(X):=f(X):=\;“Summarize {X}”. Since it is possible to obtain a summary of a summary, the arrows are composable and associative, and hence Summ is a category. We assign the internal hom to be the string descriptions of the prompts.
Let Expand be the task-category with key points for writing passages, and passages, as objects. Its arrows are prompts that take key points to passages, such as f(X):=f(X):=\;“Write a passage using the following: $\{X\}$”. Composition and associativity are symmetric to Summ: all expanded key points are new key points.
Writing a passage from keypoints is the dual of summarization: to see this, take every instruction and reverse it (e.g. $Y=\text{``Summarize }\{X\}\text{''}$ and $X=\text{``Expand \{Y\}''}$).
@@ 6.1 Prompting to Prompt
Much existing work in prompting can be construed as meta-prompting approaches, although called by different names. Below we limit ourselves to discussing strategies that fit this criterion. For an in-depth survey of prompting methods, see PromptSurvey and qiao-etal-2023-reasoning.
@@
For pretraining perhaps the clearest example is instruction pretraining (wei2022finetuned; sanh2022multitask; NEURIPS2022_b1efde53), where a generator provides synthetic tasks (prompts and context) to the model being trained. Self-instruct (selfinstruct) is an efficient-yet-noisy algorithm designed for instruction pre-training that takes in a small set of prompts and generates new prompts. LM-BFF (gao-etal-2021-making) is an approach for pre-training and fine-tuning language models via automated prompt generation. The authors show it is more effective than basic fine-tuning in various tasks.
For performance improvement, shin-etal-2020-autoprompt showed with AutoPrompt that generating prompts is more effective than fine-tuning and manual prompts in low-resource scenarios. They also showed this method lowers hallucination rates in models. More recently, zhou2023large showed that it is possible to automate some of the prompt engineering work.
Meta-prompting approaches have also found success at lowering biases (guo-etal-2022-auto), event argument extraction (ArgumentPrompt), and jailbreak detection (pryzant2023automatic). liu-etal-2022-generated showed that generating knowledge automatically for question answering (QA) improves downstream performance. For reasoning, Auto-CoT (zhang2023automatic) propose building the pathways for reasoning in chain-of-thought prompting (ChainOfThought; NEURIPS2022_8bb0d291) automatically. Closer to pure meta-prompting automation, yao2023react introduced a prompting paradigm designed for reasoning, which uses a feedback loop to generate reasoning steps. This is effective when integrating external-world grounding. LAMBADA (kazemi-etal-2023-lambada) is a backward-chaining algorithm that focuses on reasoning by extracting propositions and performing proving via chaining. Auto-GPT (AutoGPT) is a tool designed to iteratively solve problems by generating its own prompts and including a database for memory.
@@ 6.2 Category Theory
Our work is the first, to our knowledge, to formally describe applications of LLMs with category theory. However, there is a broad push from the community to apply this field to machine learning (CatMLSurvey), such as modelling gradient-based learning (GradientCat; AutoDiffEssence; Bprop), aspects of natural language processing like semantics (coecke; OpenQuantum), QA (QACat), discourse analysis (DiscourseAnalysis), and negation (NegationCoecke). These usually rely on a finer-grained approach to model language involving a functor between grammars and semantics (coecke2). The application of category theory to computation is well-known: for example, via Turing categories (LONGO1990193; COCKETT2008183). Turing categories are considerably more general than task-categories, and are more suitable to study computation as opposed to LLM behavior on specific tasks.
For the reader interested in learning more about category theory, we encourage them to read the works by FongSpivak and Spivak2 in the context of applied category theory; and MacLane, Riehl and Awodey for more in-depth content.
@@ 7 Limitations
Strictly speaking, our framework is not fully realistic. First, computable functions are partial functions, and hence so are prompts. We do not account or evaluate this. We have worked with tokens as opposed to integers. This is not a problem since for every tokenized string there exists an equivalent vector representation, and hence a correspondence with FinVect, the category of vector spaces over a finite field. Our model of isomorphism is paraphrasing: stricter definitions exist, but not without their drawbacks (Rainer).
@@
Finally, we have not modeled stochasticity, and assumed that the LLM follows the instructions. This is not necessarily the case. It is possible to model probability with a Markov category (FRITZ2023113896) such as Stoch. Markov categories are symmetric; but Prompt is not required to be.
@@ 8 Conclusion
Our work with Prompt showed that it is possible to abstract out common LLM ICL issues such as prompt sensitivity, generalizability, and user interaction; and still be able to model LLM interaction. We plan to extend our framework into a more detailed treatment of Prompt to account for stochasticity, enrichment, and the FinVect category, and link that to user preferences. Some prompting techniques, such as chain-of-thought, and its hybridization with Markov chains, such as Deepseek R1’s deepseek can also be modeled with our framework, although we leave it for future work.
@@ Appendix A Proof of Lemma 1
It then follows that any morphism in $\text{Hom}_{\text{{Task}}}(-,-)$ [^6] The notation f()f(-) is standard in category theory. It means “insert your object here.” will have (by construction) a corresponding exponential object, and hence that any natural-language description will be an element of Task.
@@ Appendix C Proof of
Although in category theory it is often sufficient to prove statements with diagrams and arguing that they commute (“proof by diagram”), in this section we provide a clearer proof of . For convenience we restate it below, and remind the reader of the properties of the hom functor. Since Prompt is right-closed, the internal hom $\hom(-,-)$, a bifunctor, mapping objects and arrows from $\text{{Prompt}}{}^{\text{op}}\times\text{{Prompt}}{}$ to Prompt. In other words, for objects $(Y,A)\in\text{{Prompt}}{}^{\text{op}}\times\text{{Prompt}}{}$ and arrows $f\colon X\rightarrow Y$, $g\colon B\rightarrow A$, for the internal hom $\text{hom}(f,g)$ the diagram below commutes
@@ D.2 Ideation
Ideation takes in a text and suggest ideas on how to improve it. The underlying task-category Idea has as objects passages, and as arrows ways to improve the (same) passage. Idea is a restricted version of Rewrite (only creative rewrites are allowed). Each object $X\in\text{{Idea}}$ has general-purpose arrows such as “Make it more concise”, “Make it longer”, and “Explain this to a 5 year old”. These three are our baseline prompts for this task. We assume that our baseline arrows belong to every exponential object $Z^{X}\in\text{{Idea}}$, but that they also have some disjoint, context-specific arrows: for example, if the context is “doing something for a long time can be boring, having hobbies and doing sports can refresh our mind”, an appropriate morphism would be “Add some details to support your claim that hobbies and sports can refresh our mind.” For our dataset we sampled 300 points from WikiText-103 (WikiText) and the TOEFL11 corpus (TOEFL). The full prompt is in Prompt 2.
@@ D.3 Creativity
The underlying task-category Creat has as objects passages, and as arrows prompts that generate said linguistic-and-semantically natural transitions between them. Same as before, we assume that the exponential objects of Creat have disjoint elements. Our baseline prompts are “Write a paragraph to connect the left text and right texts,” “Insert a passage connecting the two passages”, and the empty string (implicitly passage completion). For this task we sampled 300 points of TOEFL11, and version 3.0.0 of the DailyMail/CNN corpus (HermannKGEKSB15). The full prompt is in Prompt 3.